find an equation for the straight line through the points with coordinates
(-1,-1) and (3,-2)
please show working

Answer by mananth(16946)

the formula for the equation of line passing through (x1y1) & (x2y2)
y-y1/x-x1= y2-y1 /x2-x1
..
(x1,y1) (-1 ,-1)
(x2,y2) (3, -2)
..
y+1/x+1 = -1/4
4(y+1)= -1(x+1)
4y+4=-x-1
4y+x=-5